OpenStudy (anonymous):

\frac { x+1 }{ (x+3)(x-3) } +\frac { 4(x-3) }{ (x-3)(x+3) } +\frac { (x-1)(x-3) }{ (3-x)(x+3) } =

OpenStudy (anonymous):

\[\frac{ x+1 }{ (x+3)(x-3) } + \frac{ 4(x-3) }{ (x-3)(x+3) } +\frac{ (x-1) }{ (3-x)(x+3) }\] Is this right?

OpenStudy (anonymous):

that is the equation,yes how ever I am trying to solve it. ^_^

OpenStudy (andriod09):

the best thing to do is dothe top first, then work on the bottom. you start with this equation: \[(x+1)+4(x-3)+(x-1)\] how would you solve this?

OpenStudy (anonymous):

Ok Well the first two equations have a common denominator so they can be added. So the next step will look like \[\frac{ (x+1) + 4(x-3) }{ (x-3) (x+3) } + \frac{ (x-1) }{ (3-x) (3+x) }\] Do you agree??

OpenStudy (amistre64):

your right side has a negative implied

OpenStudy (amistre64):

3-x = -(x-3)

OpenStudy (amistre64):

just move the negative up top, and combine it with the rest